one quart of paint is exactly enough for two coats of paint on a 9foot by 10 foot wall how many quarts of paint are needed to apply one quart of paint to a 10 foot by 12 foot wall?
Found 2 solutions by mukhopadhyay, xcentaur:
Answer by mukhopadhyay(490)

One quart serves 2 coatings on 9X10 = 90 ft^2.
=> one quart serves one coating on 180 ft^2
The area of the required wall is 10X12 = 120 ft^2, and one coating is required on it.
180 ft^2 comsumes 1 quart
=> 1 ft^2 consumes 1/180 quart
=> 120 ft^2 consumes 120/180 quart = 2/3 quart

Answer by xcentaur(357)

1 quart ------------ 9(10)=90
?? quarts ------------ 10(12)=120
cross multiply
x(90)=1(120)
x=120/90=4/3=1.33 quarts of paint required


-xC